# CtRL-Sim: Reactive and Controllable Driving Agents with Offline Reinforcement Learning

### v2

@@ Appendix H CAT Simulated Data Collection and CtRL-Sim Finetuning
Method Times Preferred
CtRL-Sim 15
CAT 8
Tie 2
Table 6: The tally of votes for the pilot study. We show the breakdown of the votes for the conducted pilot study. Strong preference in terms of plausibility is shown for scenarios generated using CtRL-Sim.
@@ Appendix I Adversarial Scenario Generation User Study
We first conduct a pilot study with 5 participants to evaluate which method (CtRL-Sim vs. CAT) generates more plausible adversarial behaviours. Each participant was tasked with evaluating the same set of 25 pairs of scenarios that were randomly selected from a pool of 259 scenarios, where one scenario employed a CtRL-Sim adversarial agent and the other was a CAT adversarial agent. In both instances, the planner was a positively tilted CtRL-Sim planner. Table 6 shows the tally of the votes, where we observed strong preference for adversarial scenarios generated using CtRL-Sim.
@@
We conduct a larger study including a total of $22$ participants. We did not record any identifying information of the participants, and participants were invited on a voluntary basis. Users were presented with the user interface shown in Figure 6, where we randomize the order between both videos. Each participant responded to a total of $30$ pairs of scenarios, which took an average $12$ minutes to complete. These pairs were randomly selected from the same pool of $259$ scenarios as the pilot study. Table 7 shows the tally of this larger study. As we can see, the results of this larger study are inconclusive compared to the pilot study. We hypothesize that the framing of the question may have been ambiguous for a larger audience, and that the pairs of videos may have included less interesting scenarios (e.g, adversary rear-ending planner), resulting in many comparisons that were very difficult to judge, thus adding noise to the survey results. We plan to further study the question of plausibility in a future user study with better scenario selection.

### v3

@@ Appendix I Adversarial Scenario Generation User Study
We conduct a user study that contained a total of 24 participants to evaluate which method (CtRL-Sim vs. CAT) generates more plausible adversarial behaviours. We did not record any identifying information of the participants, and participants were invited on a voluntary basis. The user study contained a total of 271 paired scenarios, where each paired scenario consisted of two interacting agents, one controlled by a positively-tilted CtRL-Sim planner and the other controlled by an adversary. The scenario conditions of the paired scenarios are identical, except one scenario employed a CtRL-Sim adversarial agent and the other employed a CAT adversarial agent. The adversarial agent in each paired scenario is highlighted in pink so that users can distinguish the adversarial agent from the remaining agents in the scene. We do not identify the planner agent as we found this to detract users’ attention from the behaviour of the adversarial agent. Each participant was tasked with evaluating a randomly selected set of 30 paired scenarios from the pool of 271 paired scenarios. For each paired scenario, users selected in which scenario the adversarial agent was more realistic, with the option to select “Tie” if the two scenarios were sufficently similar. Users were presented with the user interface shown in Figure 6, where we randomize the order between both videos. We use the same positively tilted CtRL-Sim planner in both of the paired scenarios. Table 6 shows the tally of the user study. The results indicate convincingly that CtRL-Sim produces more plausible adversarial behaviours than CAT based on the participants’ preferences, as CtRL-Sim was preferred 54 more times than CAT over the 271 paired scenarios.
@@
Appendix M Dynamic Exponential Tilting
Tilting Goal Success Rate (%)
0 $69.0\pm 0.7$
-10 (Late) $64.1\pm 1.3$
10 $63.1\pm 2.6$
(a) Goal Tilting
Tilting Collision Rate (%)
0 $14.5\pm 1.5$
-10 (Late) $18.2\pm 2.3$
10 $28.9\pm 1.8$
(b) Vehicle-Vehicle Tilting
Tilting Offroad Rate (%)
0 $15.2\pm 0.9$
-10 (Late) $17.4\pm 1.0$
10 $21.6\pm 0.6$
(c) Vehicle-Edge Tilting
Table 8: We report the results of the CtRL-Sim FT model over 1000 test scenes applying (a) Goal Tilting, (b) Vehicle-Vehicle Tilting, and (c) Vehicle-Edge Tilting. We either apply no tilting, -10 tilting, or 0 tilting for the first half of the rollout and -10 tilting for the second half of the rollout (i.e., Late). We report the mean ± std over 5 seeds.
[…]
As expected, with late tilting, we observe a higher goal success rate and a lower collision/offroad rate compared to the model that applies negative tilting at all timesteps. Conversely, we see a lower goal success rate and a higher collision/offroad rate compared to the model with no tilting. This shows that tilting remains effective at the later timesteps. An interesting observation is that late tilting performs similarly to full tilting of the goal reward, likely because avoiding the goal can be effectively achieved within the last 4 seconds of the rollout, even if the agent was initially traveling towards the goal. On the other hand, it may be more difficult for the agent to collide with another agent or a road edge in the last 4 seconds if it hasn’t planned for that in the first 4 seconds.
Appendix N Failure Cases
Figure 10: Failure cases. Visualization of two failure cases of CtRL-Sim. The agent controlled by CtRL-Sim is depicted in teal. On the left, the agent drifts off the road, likely due to imitation drift. On the right, the agent’s behaviour is erratic and implausible due to extreme tilting values (-50).
In Figure 10, we visualize two failure cases generated by CtRL-Sim. In the scenario on the left, we control one agent using CtRL-Sim with zero tilting, indicated in green. However, the agent drifts off the road, exhibiting a failure mode likely caused by imitation drift—a common issue in imitation learning methods when rolled out over time. This drift likely results from the model’s lack of exposure to sufficient examples of agents successfully recovering their position when approaching the edge of the road. In the scenario on the right of Figure 10, we control one agent using CtRL-Sim with -50 vehicle-vehicle tilting. The agent’s behaviour is erratic and implausible, which is caused by selecting a large negative tilting value that pushes the model out of distribution. This failure mode can be mitigated by choosing tilting values that are not too extreme. We find empirically that tilting CtRL-Sim with values between -25 to 25 generally produce more plausible outputs.